OpenStudy (anonymous):

I WILL FAN/GIVE YOU A MEDAL/WRITE A TESTIMONIAL FOR YOU IF YOU HELP ME!! ANSWERS BELOW~! Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1).

OpenStudy (anonymous):

parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

A line parallel to: y = x + 5 will have same slope.

OpenStudy (anonymous):

so do i have to like use y-y1=m(x-x1) for all of them to find which one has the same slope

OpenStudy (saifoo.khan):

Yes. One equation can be found out by: y - (-1) = 1(x-(-2))

OpenStudy (anonymous):

then in case there is two with the same slope, the reciprocal will be the slope of the perpendicular line right?

OpenStudy (saifoo.khan):

The first one will be with same slope. The second one will be the reciprocal one. Which will be -1.

OpenStudy (anonymous):

thanks!